# COMPACT: COMPositional Atomic-to-Complex Visual Capability Tuning

### v1

@@ 5 Discussion
Conclusion.  In this work, we introduce COMPACT, a data recipe that systematically combines atomic visual capabilities (e.g., object recognition, spatial reasoning, shape attribution) into composite capabilities to solve complex multimodal tasks. Our experimental results show that explicit training on compositions of atomic capabilities matches the full LLaVA-665K [24] VIT in performance across benchmarks with less than 10% of its data budget. Our work presents the potential of structured compositional learning as a scalable, data-efficient pathway toward multimodal models that can solve complex, multi-capability tasks via compositional generalization.
Limitations.  Our approach faces two key limitations. First, we rely on data generated from closed-source models (i.e., Gemini), which potentially introduce their compositional limitations and biases to our dataset. Additionally, this data generation process is costly, which could pose challenges for reproducibility. To support future research, we will publicly release the data generated in this project. Second, our approach focuses on the compositionality of vision-centric capabilities. Therefore, our approach may not be optimal for addressing knowledge-intensive tasks that lie outside the scope of visual reasoning. See Appendix §A for a detailed discussion on knowledge-intensive task results.
Future Work.  We aim to extend COMPACT to accommodate higher-order compositional complexity (k $>$ 3). Currently, our data recipe only generates data up to k $=$ 3 due to the decreasing reliability of closed-source models at higher compositional complexities. Specifically, as the number of atomic capabilities increases, their integration tends to be more inconsistent, ambiguous, or erroneous. Future work could explore hierarchical composition approaches or hybrid data generation pipelines that combine multiple sources and verification steps to improve performance on higher compositional complexities. Additionally, experimenting with explicit reasoning approaches (e.g., step-by-step decomposition [33]) could further improve the model’s ability to solve complex tasks while retaining data efficiency.
Acknowledgments.  This material is based upon work supported by the National Science Foundation under Grant 2107048 and 2112562. Any opinions, findings, and conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. All experiments, data collection, and processing activities were conducted at Princeton University. Meta was involved solely in an advisory role and no experiments, data collection or processing activities were conducted on Meta infrastructure. We thank Allison Chen for helpful discussions and feedback.
Appendix
In this supplementary material, we provide additional analysis (§A). We also detail the implementation process of our compositional question generation and verification step (§B). Finally, we include visualizations (§C) that demonstrate the effectiveness of compositional tuning through comparative case studies.
@@ Appendix A Additional Analysis
Analysis of Conversation Length Distribution in LLaVA-665K Fig. 9 shows the distribution of the number of conversations per image in LLaVA-665K [24]. 93.6% of the samples fall below the 10-pair threshold. The distribution’s mean of 5.18 conversations per image ($\sigma=5.62$) shows that the data is heavily skewed towards lower values. We fix the target number of conversations per image in the compositional tuning dataset based on these findings. We ensure a fair comparison by aligning the distribution of our data with the baseline distribution.
Figure 9: Distribution of conversations per image in LLaVA-665K. The overwhelming majority of images (97.69%) have $\leq$20 conversation pairs. The average of number of conversations per image is 5.18 ($\sigma=5.62$). A small subset (2.31%) exceeds 20 conversations, which includes a sample with the maximum length of 275. Total conversations: $3,\!444,\!246$.
Figure 10: Distribution of Compositional Complexities in LLaVA-665K samples. Majority of questions (59.2%) use one atomic capability, followed by 30.9% using two.
Analysis of Limited Performance Gains on Knowledge-Intensive Tasks.  While our compositional tuning approach shows general improvements on various benchmarks, we observe more modest gains in knowledge-intensive tasks. Table 4 compares the performance of different approaches on OK-VQA, MMMU, and MMMU-Pro benchmarks. COMPACT with 32k compositional tuning data shows relatively small improvements over the random baseline: OK-VQA (50.02% vs 49.30%), MMMU (33.89% vs 32.89%), and MMMU-Pro (20.23% vs 18.15% on standard tasks, 11.91% vs 11.44% on vision tasks). Notably, training on the full LLaVA-665K [24] VIT dataset leads to limited performance improvements on MMMU (33.89%). Although knowledge-related tasks are not our main focus, this inspires future work on designing compositional tuning approaches that cover broader capabilities outside of the vision space.
Table 4: Limited Performance Improvements on Knowledge-Intensive Benchmarks. Comparison shows modest improvements over random baseline on tasks that require substantial world knowledge or domain expertise. Numbers reported in accuracy (%) and relative performance to full model (%).
@@
LLaVA-665K [24] 57.96 33.89 20.12 11.97 100%
Analysis of Compositional Complexities of Multi-Capability Benchmarks  The average compositional complexity $\bar{k}$ varies among benchmarks. We use Gemini-2.0-Flash [36] to analyze each question and identify the atomic capabilities required to give an answer (see the details of the system prompt in §B). We average these numbers for each benchmark to compute benchmark-specific $\bar{k}$ values. This reveals varying levels of compositional complexities across benchmarks: InfoVQA [27] ($\bar{k}=0.34$), SeedBench2Plus [17] ($\bar{k}=1.11$), MME [9] ($\bar{k}=1.16$), TextVQA [35] ($\bar{k}=1.19$), MMVet [43] ($\bar{k}=1.24$), CV-Bench [37] ($\bar{k}=1.33$), MMStar [6] ($\bar{k}=1.40$), and LLaVA-W [24] ($\bar{k}=3.05$).
Appendix B Additional Experiment Details
@@
We provide the system prompt for our capability analysis where we identify all the required capabilities for a given question (A). We also provide the system prompts for compositional question generation (B) and verification (C). The generation prompt includes structured guidelines to ensure that the generated multi-capability questions naturally blend different capabilities and can only be answered by checking the corresponding images. The verification prompt checks if the questions meet these guidelines and do not contain subjective interpretations or compositional flaws.
(A) System Prompt for Capability Analysis Prompt: You are an AI assistant that analyzes questions to identify the core capabilities required to answer them. Given a question, identify ALL the capabilities it requires from this list:
- spatial relationship (understanding relative positions)
- object interaction (how objects/people interact)
- object relationship (relationships between objects)
- text recognition (reading text in images)
- spatial recognition (understanding 3D space)
- action recognition (identifying actions/activities)
- object recognition (identifying objects)
- counting (counting objects/people)
- color (identifying colors)
- shape (identifying shapes)
Return ONLY a JSON array of the required capabilities, like: [“capability1”, “capability2”]
@@ Appendix C Visualizations
Qualitative Comparison.  We provide qualitative visualizations that compare the outputs from our compositionally-tuned COMPACT model and the LLaVA-665K VIT model. Examples in Fig. 11 highlight the importance of compositional tuning for handling complex multi-capability tasks ($k\geq 3$). These cases demonstrate COMPACT model’s enhanced ability to integrate multiple visual capabilities, while showing the baseline model’s difficulty with such compositionally complex queries.
Figure 11: Qualitative comparison of model outputs. Examples showing responses from our compositionally-tuned COMPACT model and LLaVA-665K [24] VIT model on complex queries that require multiple capabilities ($k\geq 3$). Our model demonstrates better integration of visual capabilities which leads to more accurate responses.
Zero-Capability Samples in LLaVA-665K We identify a subset of samples in the LLaVA-665K dataset that require no visual capabilities, which we refer to as zero-capability samples. These include general knowledge queries, subjective prompts, or requests that can be answered without inspecting the image at all. While such data may still be useful for instruction following, it does not contribute to the development of vision-centric skills. In our analysis, we find that approximately 1.1% of the questions in LLaVA-665K fall into this zero-capability category.

### v2

@@ 5 Discussion
Conclusion.  In this work, we introduce COMPACT, a data recipe that increases the complexity of the training samples by combining higher number of atomic visual capabilities (e.g., object recognition, spatial reasoning, shape attribution). Our experimental results show that training on more complex samples matches the full LLaVA-665K (llava15) VIT performance across benchmarks with less than 10% of the original data budget. Our work presents the potential of complexity-enhanced learning as a scalable, data-efficient pathway toward multimodal models that can solve multi-capability tasks.
Limitations.  Our approach faces two key limitations. First, we mainly rely on data generated from closed-source models (i.e., Gemini (team2023gemini)), which potentially introduce their compositional limitations and biases to our dataset. We provide further discussion on failure modes in Appendix §A.6. Additionally, this data generation process is costly, which could pose challenges for reproducibility. To support future research, we will publicly release the data generated in this project. Second, our approach focuses on the compositionality of vision-centric capabilities. Therefore, our approach may not be optimal for addressing knowledge-intensive tasks that lie outside the scope of visual reasoning. See Appendix §A.3 for a detailed discussion on knowledge-intensive task results.
Future Work.  We aim to extend COMPACT by investigating the intersection of image complexity and question complexity. Currently, our data recipe generates questions given fixed set of images. Specifically, certain groups of atomic visual capabilities and complexity levels might be optimal for different types of visual content present in the image. Additionally, investigating how complexity impacts learning efficiency of different learning algorithms such as reinforcement learning would be promising avenues of future research.
Acknowledgments.  This material is based upon work supported by the National Science Foundation under Grant 2107048 and 2112562, and the Solidigm AI SW. Any opinions, findings, and conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation and Solidigm Research. All experiments, data collection, and processing activities were conducted at Princeton University. Meta was involved solely in an advisory role and no experiments, data collection or processing activities were conducted on Meta infrastructure. We thank Allison Chen, Sanghyuk Chun, and Jihoon Chung for helpful discussions and feedback.
[…]
@@ Appendix A Additional Analysis
Figure 6: Comparison of capability distribution. The bar plots show the frequency of each atomic capability in LLaVA (left) and COMPACT (right) samples. In LLaVA, the distribution is notably imbalanced: object recognition and scene understanding are some of the most frequent, while shape and spatial recognition are less prevalent. In contrast, COMPACT exhibits a more balanced distribution across capability categories.
Figure 7: Distribution of implicitly assumed capabilities. The bar plot shows the frequency of atomic capabilities that are implicitly assumed during COMPACT’s compositional tuning data generation. Object recognition is most commonly assumed during question generation, most likely to generate meaningful questions with object centered images.
Figure 8: Correlation between capabilities. The heatmap shows the correlation between unique capabilities in COMPACT’s compositional tuning data. Object recognition’s correlations with other capabilities are relatively strong. Spatial capabilities are also locally correlated, as spatial questions require some understanding of the scene, depth, and relative position in practice.
Figure 9: Limited performance improvements on knowledge-intensive benchmarks. Comparison shows modest improvements over random baseline on tasks that require substantial world knowledge or domain expertise. Numbers reported in accuracy (%) and relative performance to full model (%).
@@
LLaVA-665K (llava15) 57.96 33.89 20.12 11.97 100%
Figure 10: Distribution of conversations per image in LLaVA-665K. The overwhelming majority of images (97.69%) have $\leq$20 conversation pairs. The average of number of conversations per image is 5.18 ($\sigma=5.62$). A small subset (2.31%) exceeds 20 conversations, which includes a sample with the maximum length of 275. Total conversations: 3,444,2463,\!444,\!246.
A.1 Relationship between k-value and complexity of the task
We use TextVQA (textvqa) to demonstrate that increasing the k-value makes the task progressively more challenging. Every question in this benchmark contains one key atomic capability (text recognition) in addition to potentially others. We split the TextVQA validation set according to the k-value of each question, and report the results of the LLaVA-v1.5-7b model.
k = 1 k = 2 k = 3
Density 73.6% 19.1% 2.2%
LLaVA-v1.5-7B 46.4 45.6 39.8
Table 4: Performance of LLaVA-v1.5-7b on each k-value in TextVQA. $k\geq 4$ is excluded as it contains less than 10 samples. Approximately 5% of the questions are $k=0$, which are ambiguous and do not strictly refer to the image (e.g., “Are Jordan products produced in Jordan?”).
Tab. 4 shows that the LLaVA-v1.5-7b model struggles more as other capabilities (e.g. color, spatial relation understanding) are added to text recognition. This provides further evidence that complexity measured by the k-value is related to the model’s ability to process visual information, and that the k-value characterizes the complexity of the question in a meaningful way. In short, as the k-value increases, overall task difficulty increases.
We report on the TextVQA benchmark because the results are more interpretable. Every question requires a shared atomic capability (text recognition) and has the same answer format, allowing us to isolate the effect of adding capabilities more strictly.
A.2 Breakdown of performance on MMStar
Category-specific performance.  We split the questions in MMStar (mmstar) by category (provided by the dataset) and report category-specific results on the models. Tab. 5 shows that COMPACT improves performance across most categories except math, which we exclude from our list of atomic capabilities for fair comparison with training on LLaVA-665K data. We observe particularly large improvements on science & technology (diagram and chart understanding) and instance reasoning (perception and relation understanding) categories, suggesting that COMPACT broadly generalizes to visual tasks.
MMStar Data coarse fine-grained instance logical math science and
perception perception reasoning reasoning technology
Random 49K 59.5 28.9 38.1 29.9 28.7 25.2
COMPACT 49K 61.6 28.9 41.2 31.8 28.1 27.7
Improvement +3.5% 0% +8.1% +6.4% -2.1% +9.9%
Table 5: Performance of COMPACT and random LLaVA-665K subset on each category in MMStar. COMPACT data has 33K VIT data and 16K compositional tuning data.
k-value specific performance.  We provide further analysis to explain the mechanism of improvement in COMPACT. We split the questions in MMStar by k-value and measure the performance of three models each trained on three different datasets ($k_{gen}$$k_{gen}=1$,2, and $k_{gen}$=1,2,3; Fig. 5) with progressively higher k-value distributions. Tab. 6 shows that $k_{gen}$=1,2,3 outperforms others on higher k-value ($k\geq 2$) questions. The improvement is larger when we compare $k_{gen}$=1,2,3 and $k_{gen}$=1, as opposed to $k_{gen}$=1,2,3 and $k_{gen}$=1,2. These results suggest that increasing the k-value of the training data enables the model to perform well on higher k-value questions in the test dataset.
Tab. 6 also shows that training on $k_{gen}=1,2,3$ leads to a small drop in performance on lower k-value ($k=1$) questions, indicating a trade-off between lower and higher k-value regimes. This explains why both simple and complex samples are necessary for training.
MMStar Data k = 1 k = 2 k = 3 k = 4
$k_{gen}=1,2,3$ vs $k_{gen}=1$ 49K -0.5% +3.6% +22.7% +33.5%
$k_{gen}=1,2,3$ vs $k_{gen}=1,2$ 49K -1.3% +2.6% +14.1% +9.1%
Table 6: Performance improvements on each k-value in MMStar by $k_{gen}=1,2,3$ dataset compared to $k_{gen}=1$ and $k_{gen}=1,2$.
A.3 Performance on Knowledge-Intensive Tasks
While our compositional tuning approach shows general improvements on various benchmarks, we observe more modest gains in knowledge-intensive tasks. Tab. 10 compares the performance of different approaches on OK-VQA, MMMU, and MMMU-Pro benchmarks. COMPACT with 32k compositional tuning data shows relatively small improvements over the random baseline: OK-VQA (50.02% vs 49.30%), MMMU (33.89% vs 32.89%), and MMMU-Pro (20.23% vs 18.15% on standard tasks, 11.91% vs 11.44% on vision tasks). Notably, training on the full LLaVA-665K (llava15) VIT dataset leads to limited performance improvements on MMMU (33.89%). Although knowledge-related tasks are not our main focus, this inspires future work on designing compositional tuning approaches that cover broader capabilities outside of the vision space.
A.4 Atomic capability selection
We conduct ablations on each atomic capability to analyze which capability is driving the most gains in COMPACT. For each iteration, we prepare an ablated COMPACT training dataset after we exclude each atomic capability in the data generation step (Step 2). Tab. 7 shows the average drop in relative performance across multimodal benchmarks (§4.1) when we ablate each atomic capability. The results indicate that all atomic capabilities contribute non-trivially to COMPACT’s performance.
Atomic Capability Rel. Performance Drop
w/o scene understanding -5.2%
w/o spatial relationship -4.9%
w/o text recognition -4.7%
w/o object recognition -4.0%
w/o color -3.7%
w/o counting -3.3%
w/o object interaction -3.2%
w/o spatial recognition -3.1%
w/o action recognition -2.1%
w/o shape -0.7%
Table 7: Average drop in relative performance as a result of ablating each atomic capability during 16K compositional tuning data generation in COMPACT.
Atomic Capability Selection.  The original LLaVA-665K data curation process lays out 8 of the 10 atomic capabilities in COMPACT; the two others (object interaction and shape attribution) we identify manually from inspecting the dataset. We acknowledge that alternative taxonomies are possible depending on specific training goals, and we view our 10 atomic capabilities as a practical starting point rather than a definitive set.
Non-Orthogonality of Capabilities.  We agree that perfect orthogonality is neither achievable nor necessary for our framework. Real-world visual reasoning naturally involves overlapping skills—for example, spatial reasoning often co-occurs with object recognition. The correlation analysis in Fig. 8 reveals these natural dependencies, which we view as a feature rather than a limitation. The key insight is that combining multiple capabilities, even correlated ones, requires the model to integrate more visual information, which increases information density.
Implicit Assumption of Object Recognition.  Fig. 8 shows that object recognition is frequently implicitly assumed during question generation. Object recognition serves as a foundational capability that often appears alongside other skills, similar to how reading comprehension is fundamental to many NLP tasks. Rather than undermining decomposability, this reflects the hierarchical nature of visual reasoning where basic perception enables higher-level reasoning. This foundational role is consistent with the design of vision-centric instruction tuning, where grounding responses in specific objects is essential for meaningful visual understanding.
A.5 Analysis on Data Statistics
Conversation length distribution in LLaVA-665K Fig. 10 shows the distribution of the number of conversations per image in LLaVA-665K (llava15). 93.6% of the samples fall below the 10-pair threshold. The distribution’s mean of 5.18 conversations per image ($\sigma=5.62$) shows that the data is heavily skewed towards lower values. We fix the target number of conversations per image in the compositional tuning dataset based on these findings. We ensure a fair comparison by aligning the distribution of our data with the baseline distribution.
Token-level analysis on COMPACT data.  We conduct token-level analysis comparing our COMPACT generated compositional tuning data (32k samples) with an equivalent-sized LLaVA-665K subset (32k samples). The results demonstrate that COMPACT achieves substantial token efficiency despite incorporating multiple atomic questions. Tab. 8 reveals that COMPACT uses 104.87 tokens/samples (3.36M total tokens) compared to LLaVA’s 197.42 tokens/samples (6.32M total tokens), achieving a 46.88% reduction (92.55 fewer tokens per entry). Specifically, COMPACT’s input (question) tokens average $12.83\pm 4.18$ (31% shorter than LLaVA’s $16.85\pm 25.49$), and answer tokens average $1.70\pm 0.90$ (92% shorter than LLaVA’s $21.74\pm 73.38$). The token reduction translates directly to faster training ($\sim$47% fewer tokens to process), lower storage usage (shorter sequences), and better data efficiency (more focused atomic capabilities per token).
Metric COMPACT LLaVA Difference
Input tokens (mean) $12.83\pm 4.18$ $16.85\pm 25.49$ 31% shorter
Output tokens (mean) $1.70\pm 0.90$ $21.74\pm 73.38$ 92% shorter
Tokens per Q&A turn 14.53 38.59 62% fewer
Total tokens per entry 104.87 197.42 46.88% reduction
Table 8: Token-level comparison between COMPACT and LLaVA datasets (32k entries each).
A.6 Quality Verification and Failure Mode Analysis
We conduct a systematic analysis of failure modes in our generated compositional tuning data data to ensure quality and transparency. To quantify the effectiveness of our quality control process, we perform a controlled experiment generating questions for $k_{gen}=1,2,3$ on a sample of images.
Rejection Rates and Failure Modes.  Our multi-stage filtering (described in Section 3.2 of the main paper) rejected approximately 21% of generated questions across four primary failure modes: (1) Low confidence scores (10%): questions where the VLM cannot answer confidently, often due to ambiguous phrasing or requiring information not present in the image; (2) Uninformative responses (12.5%): questions with answers like “unknown”, “not visible”, “yes”, or “no” that do not provide meaningful visual supervision (for example, “What text is visible on the wooden tray?” answered with “None”); (3) High word overlap/near-duplicates (40%): questions that share more than 60% of words with previously accepted questions for the same image (for instance, generating both “What is the color of the bench in the image?” and “What is the color of the bench located in the center of the scene?” with 70% word overlap); and (4) Capability mismatch (37.5%): questions that do not naturally integrate the specified capabilities (for example, asking “What object is present in the image without any action being performed?” for $k_{gen}=2$ with action_recognition and object_recognition, where the question only requires object recognition, as the VLM correctly identified only 1 of 2 expected capabilities). Notably, capability mismatch becomes increasingly important as $k$ increases, accounting for 0%, 50%, and 66.7% of rejections for $k_{gen}=1,2,3$ respectively. This demonstrates that our verification successfully filters questions that artificially force capability combinations.
Common Failure Patterns in Remaining Data.  Despite our rigorous filtering, some failure patterns remain in the final dataset: (1) Overly complex $k_{gen}=3$ questions: some high-complexity questions are difficult even for humans to answer, potentially due to the challenge of naturally integrating three or more capabilities in a single question; (2) VLM misidentification errors: cases where the VLM generator misidentifies image content, leading to incorrect ground-truth answers (while our verification step mitigates this issue, it does not eliminate it entirely); (3) Spatial reasoning errors: questions involving spatial relationships when objects are ambiguously positioned, leading to potential disagreement about correct answers; and (4) Attribute questions for occluded objects: color or shape questions about small or partially occluded objects where the attributes are not clearly visible. However, COMPACT achieves consistent improvements across diverse benchmarks with different evaluation protocols, suggesting these biases do not critically harm generalization. This multi-stage quality control ensures that our training data maintains both high quality and natural capability integration.
A.7 Additional Experiments with Open-source Data Generator
We evaluate COMPACT using an open-source generator (Qwen3-VL-4B-Instruct) to test whether our gains persist without proprietary generation. Tab. 9 reports these results, alongside the full LLaVA-665K baseline and our compact COMPACT recipe for direct comparison.
Table 9: Baseline comparisons (Qwen3 generator). Following Tab. 2, We report COMPACT results using an open-source generator (Qwen3-VL-4B-Instruct), along with the full LLaVA-665K (llava15) baseline and our compact COMPACT recipe for reference.
Recipe # Data InfoVQA SeedBench2Plus MME TextVQA $M$-Vet CV-Bench MMStar LLaVA-W Rel. (%)
LLaVA-665K 665K 20.80 41.72 1478.48 46.99 29.22 60.92 35.11 68.50 100.00
COMPACT (ours, Gemini) 65K 23.68 43.13 1379.94 44.37 31.74 55.28 36.13 64.50 100.18
COMPACT (ours, Qwen3) 65K 22.97 42.12 1370.76 43.07 30.78 56.38 34.04 65.70 98.31
Appendix B System Prompts
@@
System Prompts for Capability Analysis and Sample Generation.  We provide the system prompt for our capability analysis where we identify all the required capabilities for a given question (A). We also provide the system prompts for compositional question generation (B) and verification (C). The generation prompt includes structured guidelines to ensure that the generated multi-capability questions naturally blend different capabilities and can only be answered by checking the corresponding images. The verification prompt checks if the questions meet these guidelines and do not contain subjective interpretations or compositional flaws.
For conversation generation, we use Gemini-2.0-Flash (team2023gemini) with temperature 0.1, top-p 0.9, max token 1000. We have 3 runs per question on average. We use 32 parallel processes to generate 32K compositional tuning data. Each generation uses about 700 tokens on average. The total generation time is roughly 2 hours and the API cost is \$86.5.
(A) System Prompt for Capability Analysis Prompt: You are an AI assistant that analyzes questions to identify the core capabilities required to answer them. Given a question, identify ALL the capabilities it requires from this list: - spatial relationship (understanding relative positions) - object interaction (how objects/people interact) - scene understanding (understanding the background) - text recognition (reading text in images) - spatial recognition (understanding 3D space) - action recognition (identifying actions/activities) - object recognition (identifying objects) - counting (counting objects/people) - color (identifying colors) - shape (identifying shapes) Return ONLY a JSON array of the required capabilities, like: [“capability1”, “capability2”]
@@ Appendix C Visualizations
C.1 Qualitative Comparison
We provide qualitative visualizations that compare the outputs from our compositionally-tuned COMPACT model and the LLaVA-665K VIT model. Examples in Fig. 11 highlight the importance of compositional tuning for handling complex multi-capability tasks ($k\geq 3$). These cases demonstrate COMPACT model’s enhanced ability to integrate multiple visual capabilities, while showing the baseline model’s difficulty with such compositionally complex queries.
Figure 11: Qualitative comparison of model outputs. Examples showing responses from our compositionally-tuned COMPACT model and LLaVA-665K (llava15) VIT model on complex queries that require multiple capabilities ($k\geq 3$). Our model demonstrates better integration of visual capabilities which leads to more accurate responses.
C.2 Zero-Capability Samples in LLaVA-665K
We identify a subset of samples in the LLaVA-665K dataset that require no visual capabilities, which we refer to as zero-capability samples. These include general knowledge queries, subjective prompts, or requests that can be answered without inspecting the image at all. While such data may still be useful for instruction following, it does not contribute to the development of vision-centric skills. In our analysis, we find that approximately 1.1% of the questions in LLaVA-665K fall into this zero-capability category.